A |B Hughes Technology Corp. recently went public with an initial public offering in which it received a total of $92.46 million in new capital funding. The underwriter used a firm commitment offering in which the offer price was $19.05 and the underwriters spread was $1.40. Hughes also paid legal and other administrative costs of $2.85 million for the IPO. Calculate the number of shares issued through this IPO. (Enter your answer in millions. Round your answer to 2 decimal places.) New capital funding Offer price Underwriters spread Other administrative costs $ 92,460,000 19.05 1.40 2,850,000 10 12 13 14 15 16 17 18 19 Complete the following analysis. Do not hard code values in your calculations, and do not round intermediate calculations Number of shares issued
Answer #1

Let the number of shares issued be N

Offer Price = $ 19.05

Total IPO Proceeds = 19.05 x N, Underwriter Spread = $ 1.4 and Total Underwriter's Fee = 1.4 x N

Net IPO Proceeds = New Capital Funding = Total IPO Proceeds - Underwriter's Fee - Administrative Costs = $ 92.46 million

19.05 x N - 1.4 x N - 2.85 = 92.46

17.65 x N = 92.46 + 2.85 = 95.31

N = 95.31 / 17.65 ~ 5.4 million

N = 5400000